OpenStudy (anonymous):

solve for y: y+5/y+4 = y+7/y+3 +1

OpenStudy (anonymous):

you have to get all the ys on one side, and then use the quadratic equation. here's what i get 0 = y^2+10y+25, using quadratic, i get y = -5

OpenStudy (anonymous):

rosie got it and the check is easy, replace x by -5 and see if it works. do you need the steps?

OpenStudy (anonymous):

yes please ?

OpenStudy (anonymous):

is there more than one solution for this equation

OpenStudy (anonymous):

oh, here were my steps! 1. put the +1 on the same denominator \[y+5/y+4 = (y+7+y+3)/y+3\] to get \[y+5/y+4=(2y+10)/y+3\] 2. Cross multiply (y+5)*(y+3)=(2y+10)*(y+4) 3. open everything up y^2+8y+15=2y^2+18y+40 4. bring it all to one side 0=y^2+10y+25 5. Use the quadratic, or just factor since its a perfect square 0 = (y+5)^2 then isolate for y, so y=-5 hope this is a better explanation!